### v1

@@ Results
Next, to demonstrate the performance of our QRC on continuous-time data [^1] The previous time-dependent task, RF-modulation-scheme classification, concerns discrete-time data., and with a task that requires both long-term and short-term memory in the quantum reservoir, we performed the following classification task: input data assumed to have come from a source of white noise is filtered using a moving-average filter having one of three filter shapes (Gaussian, Lorentzian and inverse-power-law), and one of two window widths ($50$ ns and $600$ ns), and the task is to identify both the filter shape and window width (Fig. 4a), leading to six possible output classes. The filter functions were normalized so that the photon number distributions generated by the time-dependent displacements are identical up to the filter width. This normalization was applied to ensure that the task is not trivially solvable by just measuring the mean photon number (see Appendix F).
@@
Because all the signals used in this dataset are noise with zero mean, a linear classifier would do no better than random guessing. On the other hand, Figure 4b visually shows (using singular-value-decomposition on the output feature space) that the quantum reservoir was able to peel apart the different noise distributions. In this space, we see that the different classes are nearly all linearly-separable, with some overlap between the long-tailed but fast 50-ns inverse-power-law noise class, and the slow 600-ns Gaussian noise class. On the task of classifying over six different sources of noise, we achieved 93% accuracy (Fig. 4c) in only 2000 shots. As seen in the confusion matrix in Fig. 4d, the primary confusion at 2000 shots was distinguishing between the 50-ns inverse-power-law noise class and the 600-ns Gaussian noise class, as expected from the overlap in the SVD of the feature space.
@@
Finally, we compared the ability of our reservoir to understand long vs short correlations in input signals. For this, we deconstructed the full 6-class task into two sets of 3-class classification tasks, where each set has the same correlation length and are only distinguishable by the filter window type (see Fig. 4d and e). The class of signals with coherence length of 50 ns highlight the convenience of our input encoding scheme, i.e. feeding signals directly into the cavity mode without the need to sample the signal discretely in time. In contrast, classification of the class of signals with coherence lengths of $600$ ns require correlations of the reservoir dynamics beyond that of the measurement rate. To highlight the advantage of our scheme, we simulate the performance of a reservoir with that of a recent gate-based protocol where the input is sampled discretely in time [48]. Our simulations results, in Appendix  G, highlight the advantage of our protocol when the sampling rate of the input is slow, which can arise in experiment such as finite pulse durations and latency introduced by the FPGA classical comparison.
@@
Figure 4e looks at the participation of the different moments $\mu_{k}$ of the measurements in the classification accuracy of the 50 ns subtask (top), and the 600 ns subtask (bottom). Here, the output features are constructed by the mean $\mu_{1}$, or the off-diagonal elements of the moments $\mu_{2}$ and $\mu_{3}$ as a function of the hamming distance, allowing us to probe the contribution of the moments as a function of the locality of the correlations. For the 50 ns subtask, we see that the most important contribution is the mean, with the second order moment being the next-most important contribution, and the third-order moment being relatively unimportant. In stark contrast, the third-order moment is most important for the 600-ns subtask, suprisingly yielding nearly 90% using non-local third-order correlations alone. The ability to distinguish stochastic signals among the combined six classes demonstrate the ability of our reservoir to capture both slow and fast features of microwave signals.
To understand the role of the Hilbert space dimension on the performance of the reservoir, in Appendix  G, we simulate an extension of our quantum reservoir with multiple qubits. Our results point to an increase in classification accuracy with every additional inclusion of a qubit to the reservoir, for the same duration of input signal received.
@@ Discussion
In summary, we have experimentally realized an analog quantum reservoir computer (QRC) and demonstrated its ability to directly process microwave analog input signals without discretization, achieving high classification accuracy on three different tasks. Previous demonstrations of quantum reservoir computing have used multi-qubit, gate-based quantum reservoirs [48, 37, 10, 41, 27, 34, 21]. In contrast, we perform machine learning directly on analog signals fed into a single oscillator coupled to a transmon qubit. Intuitively, an analog (continuous-time, partially continuous-variable) quantum reservoir should be well-matched to processing microwave signals that may be continuous in time as well as amplitude. In addition to demonstrating accurate classification of microwave signals in our experiments, we also performed a direct comparison with a state-of-the-art discrete-time, gate-based QRC approach in simulation, and found that a continuous-time reservoir outperforms a discrete-time reservoir when the input signals contain temporal variations fast relative to the discretization time.
While our quantum reservoir only has two constituents (an oscillator and a qubit) we are nevertheless able to construct high-dimensional output features from the reservoir—which is essential in reservoir computing [12]—by performing multiple ($M$) projective measurements of the qubit during the dynamics between resets of the reservoir. We have proposed and demonstrated using central moments to construct output feature vectors from correlations between the measurement results. This approach has two key benefits. First, it allows one to control the dimension of the feature vector through the choice of the maximum order of correlators to include—which is important because while it is important to have high dimensionality, it is also possible for the dimensionality to be too high [^2] Two examples of disadvantages of feature-vector dimensionality being too high are: the classical post-processing and linear-layer computations may become overly costly, and the required number of shots may become too large.. This is in contrast to, for example, constructing a feature vector from the histogram of all $2^{M}$ possible bitstring outcomes of performing $M$ qubit measurements—in which case the feature vector has fixed size $2^{M}$. Second, central moments provide a natural way to extract non-trivial correlations in the measurement results, which is best explained with an example: a correlation $\braket{x_{1}x_{2}}$ may be dominated by the product $\braket{x_{1}}\braket{x_{2}}$, and we use the approach of central moments to subtract this trivial component. We performed experiments that compared the central-moment-correlators feature-vector construction with the histogram feature-vector construction and found that the former approach yielded better accuracy.
@@
For any quantum neural network, including QRC approaches, a central concern is to what extent one can achieve high accuracy on a particular task without needing an impractical number of shots [47]. Ref. [21] recently reported that certain functions—termed eigentasks—can be constructed with low error from quantum reservoirs even when the number of shots is modest, giving evidence that for some tasks, sampling noise need not be overwhelming. In our experiments, we found that it was possible to achieve high accuracy for all the tasks we attempted while needing only $10^{3}$$10^{4}$ shots (depending on the task). There is important future work to be done in exploring the tradeoffs between reservoir size (e.g., number of oscillators or qubits), number of measurements $M$ between reservoir resets, feature-vector dimension (dependent both on $M$ and the choice of order of correlators to include), and number of shots required for both training and inference. Because in our construction the feature-vector dimension can be adjusted without changing $M$, it is possible to, for example, explore the impact of feature-vector dimension and content on task accuracy while using the same number of measurements and a fixed number of shots.
Our quantum reservoir is small enough that it is easy to simulate classically, so it does not—at its present size of just one cavity and one qubit, at least—provide a quantum computational advantage. We nevertheless performed two studies to try understand what role quantumness is playing in our reservoir in achieving the classification accuracies that we experimentally observed. First, we showed that by artificially decreasing the coherence time of the qubit through injection of noise, the classification accuracy decreased. Second, we performed simulations of our QRC with a classicalized model of the quantum reservoir, in which no entanglement could be present, and found that this classicalized simulation of our QRC achieved worse accuracy than our quantum experimental results. These studies provide strong evidence that quantumness plays an important role in the operation of our quantum reservoir.
@@
With improved quantum hardware, we anticipate that it will be possible to carry out even more sophisticated tasks than what we have already demonstrated. Increasing the coherence time of the oscillator would enable us to perform many more measurements (the qubit’s coherence time is, favorably, less important in our scheme because our protocol involves repeatedly projectively measuring the qubit). While we analytically showed in Appendix H the ability of our QRC to be able to approximate any scalar function of the input signal when the signal is time-independent, provided the number of measurements $M$ performed is large enough, there remains the open theoretical question of the expressiveness of the QRC when the input signal is time-dependent. Extending the qubit-oscillator system to have multiple qubits and/or multiple oscillators would provide a larger Hilbert space and the potential for more complex dynamics and entanglement, which should in turn support more sophisticated computations.
@@
It is an open question if QRC—using the type of reservoir we considered in this paper, or any other—can, when implemented with NISQ hardware, achieve a quantum computational advantage over the best classical machine learning approaches, just as it is unclear if any quantum-machine-learning method can [8]. We did not investigate the potential for purely computational quantum advantage: our quantum reservoir is small enough to be easily classically simulable, and we did not vary its size in experiment to systematically study scaling. In the setting of processing prerecorded signals (which can be copied and replayed with negligible added noise), our single-oscillator, single-qubit QRC would offer no computational advantage over the best classical algorithms running on classical digital computers. However, our work opens up the possibility to experimentally achieve a different type of quantum advantage than a purely computational one. If one performs quantum processing on data obtained by a quantum sensor, there is the potential for an advantage that is a hybrid of being due to the advantage of quantum sensing and of quantum computing [8]. Our work suggests the feasibility of concretely realizing this kind of hybrid quantum sensing-computational advantage, where the quantum sensor is a superconducting circuit that can detect classical microwave radiation with high quantum efficiency and low noise [46, 45, 4, 3], and the processing of the received signal can happen within the same superconducting circuit as the detection occurred. Our experiments have shown that it is possible to accurately classify signals using a superconducting circuit even when there are only a few photons of signal in the superconducting circuit within any single run. Combining this with a sensitive quantum detector could lead to quantum smart sensors—quantum versions classical in-sensor processors [50]—that can reliably extract information from weak microwave signals in a way that exceeds the accuracy of any equivalent classical system.
@@
A.S. designed and carried out the hardware experiments and performed the data analysis. S.P. performed the numerical simulations of the quantum system and helped to optimize the experimental protocol with early contribution from J.K.. V.K. performed the numerical simulations of the classicalized quantum system, and performed the comparisons with classical machine-learning methods. V.F. oversaw the design and creation of the superconducting device by S.R. and others. A.S. and V.F. set up the cryogenic and microwave apparatus. S.R. calibrated the superconducting device with A.S. and V.F.. Y.C. and X.W. performed the theoretical analysis of the expressivity in Appendix H. T.O., L.G.W. and P.L.M. conceived the project, and T.O. and J.K. performed initial numerical simulations to validate the concept. A.S., S.P. and P.L.M. wrote the manuscript with input from all authors. P.L.M. supervised the project.
@@ Acknowledgements
The authors would like to thank Hakan Türeci, Shyam Shankar, Saeed Khan, Haohai Shi, William Banner, Shiyuan Ma, and Maxwell Anderson for helpful discussions and comments. The authors would also like to thank Bradley Cole, Clayton Larson, Britton Plourde, Eric Yelton, and Luojia Zhang for the fabrication of the transmon and on-chip resonator, Chris Wang for the design of the transmon, the on-chip resonator and the 3D superconducting cavity (using pyEPR [33]), and Nord Quantique for the fabrication of the 3D superconducting cavity. We gratefully acknowledge MIT Lincoln Laboratory for supplying the Josephson traveling-wave parametric amplifier (TWPA) used in our experiments. This paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0203. We gratefully acknowledge a DURIP award with AFOSR award number FA9550-22-1-0080 for equipment used in this work. The authors wish to thank NTT Research for their financial and technical support. PLM acknowledges membership in the CIFAR Quantum Information Science Program as an Azrieli Global Scholar. Y.C. and X.W. were supported by the Air Force Office of Scientific Research under Grant No. FA9550211005, NSF CCF-1942837 (CAREER), and a Sloan Research Fellowship.
Appendix Contents
[…]
[sections]0
Appendix A Summary of methods
@@
For time-independent signals, the unitary implemented in our reservoir (see Fig. 2b) can be approximated by the following set of unitaries (see Appendix C)
@@
This combination of unitaries enclose a loop in the oscillator’s phase space. The area of this closed loop, which depends on the phase of the unknown displacement $\beta$, imparts a geometric phase onto the qubit. In this work, we perform this unitary directly after a qubit measurement without reset. The action of the combined unitary on the qubit prepared in the ground or excited state, and for an arbitrary cavity state, is
@@
For general time-dependent signals, the closed loop formed by Eqs. 2-8 is broken, and the system is entangled before the measurement. While this can be hard to study analytically in the general case, we take a look at a special case of time-dependent signals, namely those of Fig. 3. Here, the signal is time-dependent up to half the duration, so that the signal is effectively two time-independent signals combined. As a result, Eqs. 4 and 6 are no longer equal, but each still a time-independent displacement, and thus, the effects of the cross-Kerr, as discussed in Appendix C, do not hinder the interpretation of the effective gate-based model. For such input signals, the state of the system just before measurement is
$|\psi\rangle=\frac{1}{2}[e^{iA_{i}}D(\beta_{i})+e^{-iA_{j}}D(\beta_{j})]\otimes|g,\mathrm{cavity}\rangle+\frac{1}{2}[e^{iA_{i}}D(\beta_{i})-e^{-iA_{j}}D(\beta_{j})]\otimes|e,\mathrm{cavity}\rangle,$ (13)
@@
The unitaries described above are followed by a qubit measurement, then a parity measurement. For time-independent signals, the qubit and cavity are disentangled at the end of the unitary, and the effect of the unitary on the cavity is just a displacement. Thus we can ignore any affects of the qubit measurement on the cavity. The state of the cavity after $M$ repeated measurements and $M$ time-independent displacements can be effectively described as
@@
where $P_{x_{n}}$ is the projector of the $n$th parity measurement $\Pi$ with measurement outcomes $x_{n}=\{+,-\}$. In Appendix H, we show that by sampling the parity measurements alone combined with the linear layer, we can realize (but not limited to) the following vector space of funtions:
@@
Here $C$ is equal to the number of classes in the data set. The largest elements of $y$ corresponds to the class that the reservoir predicts the given input data point $x$ belongs to. To train the weight matrix $W_{\mathrm{train}}$, we either use a pseudo-inverse method to minimize the mean squared error (MSE) between $W_{\mathrm{train}}x$ and $y$, or backpropagation to minimize the MSE after a softmax function. Both methods are described in more detail in Appendix E. In the main manuscript, we present results for whichever performed the best.
Appendix B Experimental setup

### v2

@@ Results
Finally, to demonstrate the performance of our QRC on continuous-time data [^1] The previous time-dependent task, RF-modulation-scheme classification, concerns discrete-time data., and with a task that requires both long-term and short-term memory in the quantum reservoir, we performed the following classification task: input data assumed to have come from a source of white noise is filtered using a moving-average filter having one of three filter shapes (Gaussian, Lorentzian and inverse-power-law), and one of two window widths ($50$ ns and $600$ ns), and the task is to identify both the filter shape and window width (Fig. 4a). The resultant dataset consisting of six classes of noisy signals was designed to probe the ability of our QRC to process high dimensional data with bandwidths larger than the cavity linewidth. Additionally, this task allowed us to probe the memory of our QRC and its ability to be sensitive to fluctuations in time, a key feature that enable temporal signal processing in QRCs [39, 30]. The filter functions were normalized so that the photon-number distributions generated by the time-dependent displacements are identical up to the filter width. This normalization was applied to ensure that the task is not trivially solvable by just measuring the mean photon number (see Appendix E).
Figure 4: Classification of filtered noise (a) We classify various fast and slow noisy signals by applying a moving average on stochastic white-noise signals. Three different filters are used for the moving averages: a Gaussian filter, a Lorentzian, and an inverse power law. For each filter, we generate stochastic analog signals based on both a 50 ns filter width, and a 600 ns filter width, with the latter being on the order of the measurement rate. Example time traces are given for the real part of an example stochastic white noise signal passed through each of the filters. We divide up these stochastic signals and send to the QRC to then distinguish the noise source. (b) Visualization of the high-dimensional output feature space using Singular Value Decomposition (SVD). Each point corresponds to a different signal over 2000 shots taken in real time (see text). (c) Classification accuracy as a function of the number of shots using third-order moments as the output feature. Our reservoir reaches 93% accuracy in about 2000 shots, corresponding to about 10 ms of the signal received. (d) Confusion matrix of the task taken at 2000 shots. (e) Participation of the mean and the off-diagonal elements of the second- and third-order moments in the classification accuracy within the subtasks of classifying different noise sources with fixed filter width. We see that for signals with long coherence times, higher-order measurement correlations are important, while for fast signals, the mean dominates in the performance contribution.
@@
Because all the signals used in this dataset are noise with zero mean, a linear classifier would do no better than random guessing. By contrast, Figure 4b visually shows (using Singular Value Decomposition (SVD) on the output feature space) that the quantum reservoir was able to peel apart the different noise distributions. On the task of classifying over six different sources of noise, we achieved 93% accuracy (Fig. 4c) in only 2000 shots. As seen in the confusion matrix in Fig. 4d, the primary confusion at 2000 shots was distinguishing between the 50-ns inverse-power-law noise class and the 600-ns Gaussian noise class, as expected from the overlap in the SVD of the feature space.
@@
Finally, we compared the ability of our reservoir to understand long vs short correlations in input signals. For this, we deconstructed the 6-class classification task into two classification subtasks, where in each subtask, the QRC learned to distinguish noisy signals generated from among three different filter window types, but with fixed window widths. The two subtasks differ by the filter window width (see Fig. 4d and e). The class of signals with coherence length of 50 ns highlights the convenience of our input encoding scheme, i.e. feeding signals directly into the oscillator mode without the need to sample the signal discretely in time. Additionally, the ability for our quantum reservoir to distinguish between signals with correlation times on the order of 50 ns demonstrates the sensitivity to signals which vary on time-scales much faster than the measurement rate. In contrast, classification of the class of signals with coherence lengths of $600$ ns requires correlations of the reservoir dynamics beyond that of the measurement rate. To highlight the advantage of our scheme, we simulated the performance of a reservoir with that of a recent gate-based protocol where the input was sampled discretely in time [49]. Our simulation results, in Appendix  F, highlight the advantage of our protocol when the sampling rate of the input is slow, which can arise in experiment such as finite pulse durations and latency introduced by the FPGA classical comparison.
@@
Figure 4e looks at the participation of the different moments $\mu_{k}$ of the measurements in the classification accuracy of the 50-ns subtask (top), and the 600-ns subtask (bottom). Here, the output features were constructed by the mean $\mu_{1}$, or the off-diagonal elements of the moments $\mu_{2}$ and $\mu_{3}$ as a function of the distance between measurements $d_{H}$, allowing us to probe the contribution of the moments as a function of the locality of the correlations. For the 50-ns subtask, we see that the most important contribution is the mean, with the second-order moment being the next-most important contribution, and the third-order moment being relatively unimportant. In stark contrast, the third-order moment is most important for the 600-ns subtask, surprisingly yielding nearly 90% classification accuracy using non-local third-order correlations alone. The ability to distinguish stochastic signals among the combined six classes demonstrates the ability of our reservoir to capture both slow and fast features of microwave signals.
@@ Discussion
In summary, we have experimentally realized an analog quantum reservoir computer (QRC) and demonstrated its ability to directly process microwave analog input signals without discretization, achieving high classification accuracy on three different tasks. Previous demonstrations of quantum reservoir computing have used multi-qubit, gate-based quantum reservoirs [49, 37, 8, 43, 26, 33, 19]. In contrast, we perform machine learning directly on analog signals fed into a single oscillator coupled to a transmon qubit. The superconducting-circuits platform not only allows us to leverage projective non-demolition (QND) non-Gaussian measurements to generate correlated output features, but is also well-matched to process microwave signals that can generally be continuous in time. In addition to demonstrating accurate classification of microwave signals in our experiments, we also performed a direct comparison with a state-of-the-art discrete-time, gate-based QRC approach in simulation, and found that a continuous-time reservoir outperforms a discrete-time reservoir when the input signals contain temporal variations fast relative to the discretization time (see Appendix F).
@@
For any quantum neural network, including QRC approaches, a central concern is to what extent one can achieve high accuracy on a particular task without needing an impractical number of shots [48]. Ref. [19] recently reported that certain functions—termed eigentasks—can be constructed with low error from quantum reservoirs even when the number of shots is modest, giving evidence that for some tasks, sampling noise need not be overwhelming. In our experiments, we found that it was possible to achieve high accuracy for all the tasks we attempted while needing only $10^{3}$$10^{4}$ shots (depending on the task). There is important future work to be done in exploring the trade-offs between reservoir size (e.g., number of oscillators or qubits), number of measurements $M$ between reservoir resets, feature-vector dimension (dependent both on $M$ and the choice of order of correlators to include), and number of shots required for both training and inference.
@@
With improved quantum hardware, we anticipate that it will be possible to carry out even more sophisticated tasks than what we have already demonstrated. Increasing the coherence time of the oscillator would enable us to perform many more measurements (the qubit’s coherence time is, favorably, less important in our scheme because our protocol involves repeatedly projectively measuring the qubit). While we analytically showed in Appendix G the ability of our QRC to be able to approximate any scalar function of the input signal when the signal is time-independent, provided the number of measurements $M$ performed is large enough, there remains the open theoretical question of the expressiveness of the QRC when the input signal is time-dependent. Generalizing our approach to spatial in addition to temporal inputs, as was explored in Ref. [42], would likely support more sophisticated computations. In Appendix F, we explore such extensions in simulations and find a marked improvement in classification accuracy.
@@
It is an open question if QRC—using the type of reservoir we considered in this paper, or any other—can, when implemented with NISQ hardware, achieve a quantum computational advantage over the best classical machine learning approaches, just as it is unclear if any quantum-machine-learning method can [7]. We did not investigate the potential for purely computational quantum advantage: our quantum reservoir is small enough to be easily classically simulable. However, our work opens up the possibility to experimentally achieve a different type of quantum advantage than a purely computational one. If one performs quantum processing on data obtained by a quantum sensor, there is the potential for an advantage that is a hybrid of being due to the advantage of quantum sensing and of quantum computing [7]. Our work suggests the feasibility of concretely realizing this kind of hybrid quantum sensing-computational advantage, where the quantum sensor is a superconducting circuit that can detect classical microwave radiation with high quantum efficiency and low noise [47, 46, 4, 3]. While the signals classified in this work originate at room temperature and are highly attenuated before reaching the device, our experiments have shown that it is possible to accurately classify signals using a superconducting circuit even when there are only a few photons of signal in the superconducting circuit within any single run. Combining this with a sensitive quantum detector could lead to quantum smart sensors—quantum versions of classical in-sensor processors [51]—that can reliably extract information from weak microwave signals in a way that exceeds the accuracy of any equivalent classical system.
@@
A.S. designed and carried out the hardware experiments and performed the data analysis. S.P. performed the numerical simulations of the quantum system and helped to optimize the experimental protocol with early contribution from J.K.. V.K. performed the numerical simulations of the classicalized quantum system, and performed the comparisons with classical machine-learning methods. V.F. oversaw the design and creation of the superconducting device by S.R. and others. A.S. and V.F. set up the cryogenic and microwave apparatus. S.R. calibrated the superconducting device with A.S. and V.F.. Y.C. and X.W. performed the theoretical analysis of the expressivity in Appendix G. T.O., L.G.W. and P.L.M. conceived the project, and T.O. and J.K. performed initial numerical simulations to validate the concept. A.S., S.P. and P.L.M. wrote the manuscript with input from all authors. P.L.M. supervised the project.
@@ Acknowledgements
The authors would like to thank Hakan Türeci, Shyam Shankar, Saeed A. Khan, Haohai Shi, William P. Banner, Shi-Yuan Ma, and Maxwell Anderson for helpful discussions and comments. The authors would also like to thank Bradley Cole, Clayton Larson, Britton Plourde, Eric Yelton, and Luojia Zhang for the fabrication of the transmon and on-chip resonator, Chris Wang for the design of the transmon, the on-chip resonator and the 3D superconducting cavity (using pyEPR [32]), and Nord Quantique for the fabrication of the 3D superconducting cavity. We gratefully acknowledge MIT Lincoln Laboratory for supplying the Josephson traveling-wave parametric amplifier (TWPA) used in our experiments. This paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0203. We gratefully acknowledge a DURIP award with AFOSR award number FA9550-22-1-0080 for equipment used in this work. The authors wish to thank NTT Research for their financial and technical support. PLM acknowledges membership in the CIFAR Quantum Information Science Program as an Azrieli Global Scholar. Y.C. and X.W. were supported by the Air Force Office of Scientific Research under Grant No. FA9550211005, NSF CCF-1942837 (CAREER), and a Sloan Research Fellowship.
Methods
@@
For time-independent signals, the unitary implemented in our reservoir (see Fig. 2b) can be approximated by the following set of unitaries (see Appendix B)
@@
This combination of unitaries encloses a loop in the oscillator’s phase space. The area of this closed loop, which depends on the phase of the unknown displacement $\beta$, imparts a geometric phase onto the qubit. In this work, we perform this unitary directly after a qubit measurement without reset. The action of the combined unitary on the qubit prepared in the ground or excited state, and for an arbitrary oscillator state, is
@@
For general time-dependent signals, the closed loop formed by Eqs. 2-8 is broken, and the system is entangled before the measurement. While this can be hard to study analytically in the general case, we take a look at a special case of time-dependent signals, namely those of Fig. 3. Here, the signal is time-dependent up to half the duration, so that the signal is effectively two time-independent signals combined. As a result, Eqs. 4 and 6 are no longer equal, but each still a time-independent displacement, and thus, the effects of the cross-Kerr, as discussed in Appendix B, do not hinder the interpretation of the effective gate-based model. For such input signals, the state of the system just before measurement is
$|\psi\rangle=\frac{1}{2}[e^{iA_{i}}D(\beta_{i})+e^{-iA_{j}}D(\beta_{j})]|g,\mathrm{cavity}\rangle+\frac{1}{2}[e^{iA_{i}}D(\beta_{i})-e^{-iA_{j}}D(\beta_{j})]|e,\mathrm{cavity}\rangle,$ (13)
@@
The unitaries described above are followed by a qubit measurement, then a parity measurement. For time-independent signals, the qubit and oscillator are disentangled at the end of the unitary, and the effect of the unitary on the oscillator is just a displacement. Thus we can ignore any affects of the qubit measurement on the oscillator. The state of the oscillator after $M$ repeated measurements and $M$ time-independent displacements can be effectively described as
@@
where $P_{x_{n}}$ is the projector of the $n$th parity measurement $\Pi$ with measurement outcomes $x_{n}=\{+,-\}$. In Appendix G, we show that by sampling the parity measurements alone combined with the linear layer, we can realize (but not limited to) the following vector space of funtions:
@@
Here $C$ is equal to the number of classes in the data set. The largest elements of $y$ corresponds to the class that the reservoir predicts the given input data point $x$ belongs to. To train the weight matrix $W_{\mathrm{train}}$, we either use a pseudo-inverse method to minimize the mean squared error (MSE) between $W_{\mathrm{train}}x$ and $y$, or backpropagation to minimize the MSE after a softmax function. Both methods are described in more detail in Appendix D. In the main manuscript, we present results for whichever performed the best.
Appendix Contents
[…]
[sections]0
Appendix A Experimental setup